# of the a an and or to from in as about but than then so through

### v1

@@ 3.2. Project Selection
We collect metadata (step 2) for the sampled repositories using GitHub’s Rest API. The metadata includes the number of lines of code, repository creation date, last push date, number of stars, number of issues, and the primary programming language of the repository. As noted by Kalliamvakou et al. (Kalliamvakou et al., 2014), one of the challenges in mining GitHub is that a large portion of repositories are not software projects. In addition, most of them are inactive and contain very few commits. Our target population consists of nontrivial or ‘non-toy’ software projects, but defining what is nontrivial is difficult and subject to the authors’ interpretation. As Maj et al. (Maj et al., 2024) suggest, it is easier to discard projects we know are not part of the population than to define what is in it.
We exclude a project from our dataset if any of the following hold (step 3): it contains less than 50 lines of code, it is aged less than 2 months, or if GitHub does not assign a primary language to it. We define the age of a project as the time between its creation date and the date of the last push to it. An alternative or complement to the age is the number of commits, though in practice querying the number of commits using the API is expensive. Discarding projects with no primary language is a way to filter out templates and projects containing only documentation or binary files. We do not filter based on social engagement, i.e., stars or issue counts, as it is unclear whether analyzing only these repositories would generalize (Maj et al., 2024; Borges and Tulio Valente, 2018). While the number of contributors could be used as a filter, we choose not to as single-maintainer projects could still reasonably be part of the population.
@@ 3.3. Language Filtering
For each programming language recognized by GitHub, we determine whether it is statically typed (based on its documentation) and, if so, include it in our list of target languages. We exclude languages that do not have primitive floating-point types (e.g., Coq, Agda, or Dafny). We also exclude languages where we could not find the language reference (e.g., Mirah or ActionScript). This yields the list of 51 statically typed programming languages in Fig.˜2. We discard repositories that do not contain code written in at least one of them (step 5).
@@ 3.4. Downloading and Keyword Filtering
We download (step 6) the content of the filtered repositories using the hash of the latest commit on the main branch we retrieve in the previous step. We exclude files not written in a statically typed language. We determine programming language using file extensions, as repositories do not contain metadata for mapping language to files. We then compute the number of lines of code and the number of words.
We aim to retain files with functions operating on floating-point values. The most precise approach to identify these would be to infer the type of each individual operation, i.e., run a type checker or type inference. This approach, however, would require special treatment and significant effort for each individual programming language. We instead choose an approximate, but more scalable, approach, and aim to identify functions with floating-point operations based on keywords used in the source code via regular expressions. Because the programming languages we are dealing with are all statically typed, we know that function signatures often contain the types of the parameters and return values. Hence, we search for floating-point types in the files using a regular expression. A problem encountered by this approach is the presence of type aliases, structures or classes that represent or contain floating-point values and that might be defined in other files. To mitigate this issue, we also look at other keywords related to floating-point computations, such as transcendental functions, rounding functions, and special values like NaN or infinity. Looking at the presence of transcendental functions is also interesting per se, as they are a feature that several floating-point verification tools aim to support.
For each of the 51 languages, we manually inspect the official language reference to identify floating-point types and special values, and we examine the corresponding standard math library to identify numerical functions. We group keywords into three categories:
  • Floating point types: keywords used to declare floating point types, e.g., float, double, real, float32, float64.
@@
We upload the JSON files containing the keyword definitions for each programming language as supplementary material.
The search is case insensitive using boundaries such as spaces, punctuation or brackets to surround the words. While choosing keywords, we tried to be as exhaustive as possible while avoiding ambiguous keywords. For example, keywords like abs or pow are valid floating-point functions but they are often used in integer computations as well, and could lead to false positives. While some keywords like cos, sin or float are common to most languages, others can be specific to a given programming language and ambiguous in others. This is the case for the keyword number in Typescript, which may denote floating-points or integers, or real in Pascal or Fortran. For each file, we count the number of keyword occurrences per category, and discard the file if it does not contain any keyword from neither of the three categories.
@@ 3.5. De-Duplication
Removing duplicates is a crucial step to avoid measurement bias, as prior work shows that GitHub code contains considerable amounts of code duplication (Lopes et al., 2017). We eliminate files that are identical up to token reordering across repositories (step 7). For this step and the following ones, we abandon the notion of repositories and treat files as standalone entities within a global pool. We represent each file as a bag of words, ignoring whitespace, delimiters, special characters, and token order. This representation enables the detection of token-level duplicates that differ only by whitespace changes or reordered declarations. We compute the BLAKE3 hash of each bag-of-words representation and retain a single instance for each unique hash.
@@ 3.6. Parsing and Function Extraction
Once only deduplicated files containing numerical keywords are left, we parse them to extract individual functions (step 8). We discard files with syntax errors, construct parse trees for valid files, and remove comments during preprocessing. We remove comments because they are not directly relevant to our analysis, may be outdated, and can bias results. In particular, terms such as double may appear in natural language text with meanings unrelated to floating-point types, thereby increasing the risk of false positives. Comments are clearly valuable in their own right and deserve special care, but we leave their analysis for future work.
@@
Similar to comments, we exclude string literals from the keyword search, since keywords occurring in free text are more likely to produce false positives. If a function contains at least one numerical keyword, we extract its code and store it in a separate file (step 9).
We parse the files using tree-sitter (Tre, 2025), a language agnostic parsing library that already comes with grammars for most used programming languages. Our implementation currently supports C, C++, Java, TypeScript, C#, Go, Fortran, which together account for more than 92 % of the functions collected. Supporting additional languages requires only defining the parse nodes of interest, which makes our implementation readily reusable and extendable in future code studies. While other statistics, such as the number of floating-point arithmetic operations, would have been relevant for the study, these rely on type information which is beyond tree-sitter’s capacity [^4] Tree-sitter preserves type information that is explicitly declared in source code, but does not do any type-checking, propagation or inference.; i.e., they would require programming language-specific analyses.
@@ 4. Experimental Setup
Our approach is implemented in Rust and designed as a general-purpose framework for large-scale code studies, with a focus on reproducibility and reusability. Data collection can be interrupted and resumed at any time without compromising reproducibility. Inputs are randomly shuffled before any pipeline step to enable subsampling. All random seeds are fixed and can be user-defined. During language collection, the hash of the latest commit on the main branch is stored, ensuring that repository contents remain consistent across re-runs, provided it is still reachable.
The command-line interface provides users with flexible control over the analysis process. For example, users can choose to analyze local repositories instead of downloading data from GitHub. They can also specify which file extensions or programming language to retain in the dataset, and adapt the set of keywords depending on the programming language. Multiple configuration files can be supplied simultaneously, enabling parallel analyses with different settings. When downloading data, users have the option to either delete files that do not pass filtering criteria or keep them for future use. Every pipeline step can be run independently as a separate subcommand in the command line interface. Each of them can be skipped or performed manually by editing the corresponding CSV file independently without affecting the others, and new steps can be added without disrupting existing functionality. The code is fully documented, tested and will be released as open source. We conduct the data analysis using Jupyter Notebooks.
@@
We run our pipeline on an Ubuntu 24.04 machine with 128 GB RAM, 12 CPU cores at 4.95 GHz, using Rust 1.85.0 and Python 3.9.6. We ran the different pipeline steps in an interleaved fashion. Since the exact running time of each pipeline step depends directly on the number of sampled projects, we report here approximate durations that are aimed to mostly illustrate the relative differences between the pipeline steps. Steps that rely on the GitHub API are rate-limited to 5000 requests per hour per token, hence the durations vary depending on the number of available tokens. For our experiments, we used 8 tokens.
Sampling IDs took approximately a week, metadata collection took 2 months, language collection took a few days, downloading the files took 3 days, removing duplicates took 10 minutes and parsing took 40 minutes. We collected the data between October 2024 and March 2025.
@@
We first provide statistics about steps 1 - 5 of our pipeline (see Fig.˜1), which are largely not specific to floating-point arithmetic, before discussing the floating-point specific results in Section˜5.
@@
At the time of metadata collection, 4.6 % of the projects were no longer reachable, having been deleted or made private. Metadata for the remaining reachable projects is summarized in the first column of Table˜1. For each statistic, the table reports the minimum, median, and maximum values, as well as the first (Q1) and third (Q3) quartiles.
@@
Table 1. Distribution of project metadata before and after metadata filtering, and after keyword filtering
[t] Metadata All “Interesting” Contain floating point keywords Min Q1 Median Q3 Max Min Q1 Median Q3 Max Min Q1 Median Q3 Max Age –1949 d 1 min 36 min 11 d 6104 d 60 d 125 d 316 d 925 d 6104 d 60 d 118 d 288 d 840 d 6089 d Creation YY/MM 70/01 20/02 22/02 23/08 25/01 08/02 19/05 20/10 22/11 24/10 08/03 19/01 20/12 22/12 23/10 Size (in LOC) 0 2 56 928 109 M 50 322 1836 9759 105 M 50 372 2111 14460 42 M # forks >> 0 7 % 20 % 24 % # forks among >> 0 1 1 1 3 144 k 1 1 2 5 144 k 1 1 2 6 27 k # issues >> 0 8 % 31 % 30 % # issues among >> 0 1 1 2 7 70 k 1 1 5 14 70 k 1 1 4 13 15 k Q1 and Q3 denote the first and third quartiles. The gray columns are discussed in Section˜5.
@@
In summary, the dataset is skewed towards small, recent and inactive projects. More than 75 % of the projects are younger than two weeks, and 85 % younger than two months. Empty repositories account for 19 % of the dataset, while 49 % contain less than 50 lines of code. Additionally, 12 % of the projects are non-empty but have no primary language assigned by GitHub, confirming the need for filtering. A small fraction (0.6 %) of the projects have a negative age, a consequence of GitHub relying on the local system time of the machine pushing the commit, which may be incorrectly configured.
@@
After removing projects with fewer than 50 lines of code, that were less than 2 months old, or contained no source code (step 3), we were left with 1 238 7411\,238\,741 projects (10.81 % of the initially collected metadata). We refer to these as the “interesting” projects in the second column of Table˜1 which shows their metadata distribution. While the number of projects created on GitHub has grown exponentially, with the majority in the last five years, this trend is less pronounced in the filtered dataset. The median creation year of retained projects is two years earlier than that of the unfiltered dataset.
@@
The dataset is dominated by projects written primarily in non-statically typed languages. According to GitHub’s language classification, JavaScript accounts for the largest share at 22 %, followed by HTML (12 %) + CSS (4 %) and Python (10 %) + Jupyter Notebook (5 %). These proportions are similar before and after filtering, except for Python, which drops from 17 % to 10 %. Among statically typed languages, the most common are Java (9 %), TypeScript (7 %) and C++ (4 %). After collecting the list of all languages used in each repository (step 4), and removing projects that do not include any statically typed language (step 5), we are left with 465 050465\,050 repositories, corresponding to 38 % of the filtered dataset that we use to answer our research questions.
5. Results
@@
  • RQ1. How prevalent is floating-point arithmetic in open-source statically typed code on GitHub? We measure prevalence at both the project and file levels. First, we identify how many repositories and files contain floating-point code. At the file level, we further investigate whether floating-point operations are distributed throughout the file or concentrated in a few functions, and provide the complete answer to RQ1 on page 5.3.
  • RQ2. What are the characteristics of floating-point code in terms of programming language constructs and size? We analyze the structure of floating-point functions in our dataset to identify commonly used programming language constructs, such as loops, branches, functions calls, and their combinations.
@@
5.1. Project-based Analysis
At the time of downloading (step 6 in Fig.˜1), 2.2 % of the projects were unreachable, leaving 454 865454\,865 accessible projects. From these, we now consider only those projects that contain floating-point keywords. Metadata for these projects is summarized in the third (grey) column of Table˜1. While not the same, the metadata is quite similar to the one of the “interesting” projects.
Table˜2 shows the breakdown of the different floating-point keyword categories by projects and files. 67 % of the projects contain at least one floating-point keyword. Most often these are floating-point types, which appear in 64 % of the projects. Transcendental functions and keywords such as NaN appear less frequently.
For each project containing at least one floating-point keyword from a category, we compute the proportion of statically typed files containing at least one keyword from that category. The median and quartiles of these proportions are shown in the second column of the table. What we observe is that when floating-point related keywords are present in a project, they tend to be concentrated in a small number of files. In half of the projects that contain floating-point types, fewer than 17 % of the files include such keywords. This proportion is even lower for transcendental functions and miscellaneous keywords.
We note that we compute these statistics before deduplication. This means that these statistics will likely contain some library code for projects but not for all, depending on which code has been committed to the repository. We also compute these statistics on the file level after deduplication in Fig.˜3.
@@ 5.1. Project-based Analysis
Floating point types 64 % 7.1 % 17 % 35 %
@@
Other keywords 21 % 1.3 % 3.0 % 7.1 %
Any 66 % 7.4 % 17 % 36 %

### v2

@@ 3.2. Project Selection
We collect metadata (step 2) for the sampled repositories using GitHub’s REST API. The metadata includes the number of lines of code, repository creation date, last push date, number of stars, number of issues, and the primary programming language of the repository. As noted by Kalliamvakou et al. (kalliamvakouPromisesPerilsMining2014), one of the challenges in mining GitHub is that a large portion of repositories are not software projects. In addition, most of them are inactive and contain very few commits. Our target population consists of nontrivial or “non-toy” software projects, but defining what is nontrivial is difficult and subject to the authors’ interpretation. As Maj et al. (majFaultOurStars2024) suggest, it is easier to discard projects we know are not part of the population than to define what is in it.
We exclude a project from our dataset if any of the following hold (step 3):
  • it contains less than 50 lines of code,
  • GitHub does not assign a primary language to it (these are typically templates, or projects containing only documentation or binary files),
  • the creation date and last activity date are less than two months apart (so the "age" of the project is less than two months).
The rationale for using age as an exclusion criterion is to avoid including repositories in the population that were e.g., used for student assignments. An alternative or complement to the age is the number of commits, though in practice querying the number of commits using the API is expensive.
We do not filter based on social engagement, i.e., stars or issue counts, as it is unclear whether analysing only these repositories would generalise (majFaultOurStars2024; borgesWhatsGitHubStar2018). While the number of contributors could be used as a filter, we choose not to as single-maintainer projects could still reasonably be part of the population.
@@ 3.3. Language Filtering
For each programming language recognised by GitHub, we determine whether it is statically typed (based on its documentation) and, if so, include it in our list of target languages. We exclude languages that do not have primitive floating-point types (e.g., Coq, Agda, or Dafny). We also exclude languages where we could not find the language reference (e.g., Mirah or ActionScript). This yields the list of 51 statically typed programming languages in Figure˜2. We discard repositories that do not contain code written in at least one of them (step 5).
@@
List of statically typed programming languages analyzed, including C, C++, Java, Rust, Go, Swift, Scala, Kotlin, and Fortran.
@@ 3.4. Downloading and Keyword Filtering
We download (step 6) the content of the filtered repositories using the hash of the latest commit on the main branch we retrieved in the previous step. We exclude files not written in a statically typed language. We determine programming language using file extensions, as repositories do not contain metadata for mapping language to files. We then compute the number of lines of code and the number of words.
We aim to retain files with functions operating on floating-point values. The most precise approach to identify these would be to infer the type of each individual operation, i.e., run a type checker or type inference. This approach, however, would require special treatment and significant effort for each individual programming language. We instead choose an approximate, but more scalable, approach, and aim to identify functions with floating-point operations based on keywords used in the source code via regular expressions. Because the programming languages we are dealing with are all statically typed, we know that function signatures often contain the types of the parameters and return values. Hence, we search for floating-point types in the files using a regular expression. A problem encountered by this approach is the presence of type aliases, structures or classes that represent or contain floating-point values and that might be defined in other files. To mitigate this issue, we also look at other keywords related to floating-point computations, such as transcendental functions, rounding functions, and special values like NaN or infinity. Looking at the presence of transcendental functions is also interesting per se, as they are a feature that several floating-point verification tools aim to support. Finally, we examine the use of keywords related to arbitrary-precision floating-point computations to assess how frequently such features are used in practice.
For each of the 51 languages, we manually inspect the official language reference to identify floating-point types and special values, and we examine the corresponding standard math library to identify numerical functions. We group keywords into four categories:
  • Primitive floating point types: keywords used to declare standard floating point types, e.g., float, double, real, float32, float64.
@@
  • Arbitrary precision floating-point functions and types, e.g., mpfr_t, bigfloat, apfloat.
The JSON files containing the keyword definitions for each programming language are available along with our function dataset.
The search is case insensitive using boundaries such as spaces, punctuation or brackets to surround the words. While choosing keywords, we tried to be as exhaustive as possible while avoiding ambiguous keywords. For example, keywords like abs or pow are valid floating-point functions but they are often used in integer computations as well, and could lead to false positives. While some keywords like cos, sin or float are common to most languages, others can be specific to a given programming language and ambiguous in others. This is the case for the keyword number in Typescript, which may denote floating-points or integers, or real in Pascal or Fortran. For each file, we count the number of keyword occurrences per category, and discard the file if it does not contain any keyword from neither of the four categories.
@@ 3.5. De-Duplication
Removing duplicates is a crucial step to avoid measurement bias, as prior work shows that GitHub code contains considerable amounts of code duplication (lopesDejaVuMapCode2017). We eliminate files that are identical up to token reordering across repositories (step 7). For this step and the following ones, we abandon the notion of repositories and treat files as standalone entities within a global pool. We represent each file as a bag-of-words, ignoring whitespace, delimiters, special characters, and token order. This representation enables the detection of token-level duplicates that differ only by whitespace changes or reordered declarations. We compute the BLAKE3 hash of each bag-of-words representation and retain a single instance for each unique hash.
@@ 3.6. Parsing and Function Extraction
Once only deduplicated files containing numerical keywords are left, we parse them to extract individual functions (step 8). We remove comments because they are not directly relevant to our analysis, may be outdated, and can bias results. In particular, terms such as double may appear in natural language text with meanings unrelated to floating-point types, thereby increasing the risk of false positives. Comments are clearly valuable in their own right and deserve special care, but we leave their analysis for future work.
@@
We also record the function’s location in the source file and whether a parsing error occurred, including its position when applicable. Similar to comments, we exclude string literals from the keyword search, since keywords occurring in free text are more likely to produce false positives. If a function contains at least one numerical keyword, we extract its code and store it in a separate file (step 9).
We parse the files using tree-sitter (Treesitter2025), a language agnostic parsing library that already comes with grammars for the most used programming languages. Our implementation currently supports C, C++, Java, TypeScript, C# and Go, which together account for more than 92 % of the functions collected. Supporting additional languages requires only defining the parse nodes of interest, which makes our implementation readily reusable and extendable in future code studies. While other statistics, such as the number of floating-point arithmetic operations, would have been relevant for the study, these rely on type information, which is beyond tree-sitter’s capacity [^7] Tree-sitter preserves type information that is explicitly declared in source code, but does not do any type-checking, propagation or inference.; i.e., they would require programming language-specific analyses.
@@ 4. Experimental Setup
Our approach is implemented in Rust and designed as a general-purpose framework, called Scyros, for large-scale code studies, with a focus on reproducibility and reusability. Data collection can be interrupted and resumed at any time without compromising reproducibility. Inputs are randomly shuffled before any pipeline step to enable subsampling. All random seeds are fixed and can be user-defined. During language collection, the hash of the latest commit on the main branch is stored, ensuring that repository contents remain consistent across re-runs, provided it is still reachable and history was not rewritten.
Scyros can be straightforwardly adapted in a number of ways for different analyses. For example, users can choose to analyse local repositories instead of downloading data from GitHub. They can also specify which file extensions or programming language to retain in the dataset, and adapt the set of keywords for each programming language depending on the target code features of interest. Multiple configuration files can be supplied simultaneously, enabling parallel analyses with different settings. When downloading data, users have the option to either delete files that do not pass filtering criteria or keep them for future use. Each pipeline step produces a CSV file that serves as the input for the next step. All steps can be executed independently as separate subcommands via the command-line interface. Individual steps can be skipped or performed manually by editing their corresponding CSV files without affecting other stages of the pipeline. Furthermore, additional steps can be introduced without disrupting existing functionality. The code is fully documented, tested, and released as open-source software under the Apache 2.0 license. We conduct the data analysis in Python.
@@
We run our pipeline on an Ubuntu 24.04 machine with 128 GB RAM, 12 CPU cores at 4.95 GHz, using Rust 1.85.0 and Python 3.13.9. We ran the different pipeline steps in an interleaved fashion. Since the exact running time of each pipeline step depends directly on the number of sampled projects, we report here approximate durations that are aimed to mostly illustrate the relative differences between the pipeline steps. Steps that rely on the GitHub API are rate-limited to 5000 requests per hour per token, hence the durations vary depending on the number of available tokens. For our experiments, we used 8 tokens.
Sampling IDs took approximately a week, metadata collection took 2 months, language collection took a few days, downloading the files took 3 days, removing duplicates took 10 minutes and parsing took 60 minutes. We collected the data between October 2024 and March 2025.
@@
We first provide statistics about steps 1–5 of our pipeline (see Figure˜1), which are largely not specific to floating-point arithmetic, before discussing the floating-point specific results in Section˜5.
@@
At the time of metadata collection, 4.6 % of the projects were no longer reachable, having been deleted or made private. Metadata for the remaining reachable projects is summarised in the first row of Table˜1. For each statistic, the table reports the minimum, median, and maximum values, as well as the first (Q1) and third (Q3) quartiles.
@@
Table 1. Distribution of project metadata before and after metadata filtering, and after keyword filtering. Q1 and Q3 denote the first and third quartiles. The gray row is discussed in Section˜5.
Metadata Age Creation Size # forks # issues
(YY/$M$) (in LOC) > 0 among > 0 > 0 among > 0
All Min –19459 d 70/01 0 6 % 1 6 % 1
Q1 1 min 20/02 2 1 1
Median 36 min 22/02 56 1 2
Q3 11 d 23/08 928 3 7
Max 6104 d 25/01 109 M 144 k 70 k
Interesting Min 60 d 08/02 50 20 % 1 31 % 1
Q1 125 d 19/05 322 1 1
Median 316 d 20/10 1836 2 5
Q3 925 d 22/11 9759 5 14
Max 6104 d 24/10 105 M 144 k 70 k
With FP
keywords
Min 60 d 08/03 50 24 % 1 30 % 1
Q1 119 d 19/01 372 1 1
Median 290 d 20/12 2102 2 4
Q3 844 d 22/12 14422 6 13
Max 6090 d 24/10 42 M 27 k 15 k
@@
In summary, the dataset is skewed towards small, recent and inactive projects. More than 75 % of the projects are younger than two weeks, and 85 % younger than two months. Empty repositories account for 19 % of the dataset, while 49 % contain less than 50 lines of code. Additionally, 12 % of the projects are non-empty but have no primary language assigned by GitHub, confirming the need for filtering. A small fraction (0.61 %) of the projects have a negative age, a consequence of GitHub relying on the local system time of the machine pushing the commit, which may be incorrectly configured.
@@
After removing projects with fewer than 50 lines of code, that were less than 2 months old, or contained no source code (step 3), we were left with 1 238 7411\,238\,741 projects (11 % of the initially collected metadata). We refer to these as the “interesting” projects in the second row of Table˜1 which shows their metadata distribution. While the number of projects created on GitHub has grown exponentially, with the majority in the last five years, this trend is less pronounced in the filtered dataset. The median creation year of retained projects is two years earlier than that of the unfiltered dataset.
@@
The dataset is dominated by projects written primarily in non-statically typed languages. According to GitHub’s language classification, JavaScript accounts for the largest share at 22 %, followed by HTML (12 %) + CSS (4 %) and Python (10 %) + Jupyter Notebook (5 %). These proportions are similar before and after filtering, except for HTML, which drops from 17 % to 12 %. Among statically typed languages, the most common are Java (9 %), TypeScript (7 %) and C++ (4 %). After collecting the list of all languages used in each repository (step 4), and removing projects that do not include any statically typed language (step 5), we are left with 465 050465\,050 repositories, corresponding to 38 % of the filtered dataset. This dataset forms the basis for answering our research questions.
5. Quantitative Analysis of Floating-Point Usage
@@
  • RQ1. How prevalent is floating-point arithmetic in open-source statically typed code on GitHub? The aim is to confirm (or disprove) the conventional hypothesis, thus (in)validating the motivation for much past and future research.
    We measure prevalence at both the project and file levels. First, we identify how many repositories and files contain floating-point code. At the file level, we further investigate whether floating-point operations are distributed throughout the file or concentrated in a few functions. We provide the complete answer to RQ1 on page 5.3.
  • RQ2. What are the characteristics of floating-point code in terms of programming language constructs and size? The aim is to characterise "real-world" floating-point code to inform future floating-point tool development and benchmarking towards practical relevance.
    We analyse the structure of floating-point functions in our dataset to identify commonly used programming language constructs, such as loops, branches, functions calls, and their combinations.
@@
Primitive floating-point types 64 % 7.1 % 17 % 35 %
@@
Miscellaneous keywords 21 % 1.4 % 3.1 % 7.1 %
Arbitrary-precision keywords 1.1 % 0.13 % 1.4 % 3.6 %
Any 66 % 7.5 % 18 % 36 %
5.1. Project-Based Analysis
At the time of downloading (step 6 in Figure˜1), 3.8 % of the projects were unreachable, leaving 447 209447\,209 accessible projects. From these, we now consider only those projects that contain floating-point keywords. Metadata for these projects is summarised in the third (grey) row of Table˜1. While not the same, the metadata is quite similar to the one of the “interesting” projects.
Table˜2 shows the breakdown of the different floating-point keyword categories by projects and files. Floating-point keywords appears in 66 % of the projects. Most often these are floating-point types, which appear in 64 % of the projects. Transcendental functions and keywords such as NaN appear less frequently explicitly in the code, while arbitrary-precision keywords appear very rarely (1.1 % of the projects).
For each project containing at least one floating-point keyword from a category, we compute the proportion of statically typed files containing at least one keyword from that category. The median and quartiles of these proportions are shown in the second column of the table. What we observe is that when floating-point related keywords are present in a project, they tend to be concentrated in a small number of files. In half of the projects that contain floating-point types, fewer than 17 % of the files include such keywords. This proportion is even lower for transcendental functions as well as miscellaneous and arbitrary-precision keywords.
We note that we compute these statistics before deduplication. This means that these statistics will likely contain some library code for projects but not for all, depending on which code has been committed to the repository.